The sum of three numbers is 44. The difference between the largest and the smallest is 14, and the largest is equal to the sum of the two smaller numbers. Find the numbers.
Answer by checkley77(12569)

x+y+z=44
z-x=14
z=x+y
x+y+(x+y)=44
2x+2y=44
x+y+z=44 ( multiply this equation by -2 *& add it to the previous equation
-2x-2y-2z=-88
2x+2y=44
-----------------------------
-2z=-44
z=-44/-2
z=-44/-2
z=22 ans.
22-x=14
-x=14-22
-x=-8
x=8 ans.
22=8+y
y=22-8
y=14 ans
Proof:
x+y+z=44
8+14+22=44
44=44